Graffiti Advertising, Inc., reported the following financial statements for the last two years 2016 Income Statement Sales Costs of goods sold Selling and administrative Depreciation EBIT Interest EBT Taxes Net income Dividends Addition to retained earnings $565,700 274,020 124,732 54,575 $112,373 19,318 $ 93,055 37,222 $ 55,833 $ 9,700 $ 46,133 GRAFFITI ADVERTISING, INC Balance Sheet as of December 31, 2015 Cash Accounts receivable Inventory Current assets $ 13,330 Accounts payable $ 9,503 14,507 $ 24,010 18,993 $ 46,118 S344,456 $390,574 Notes payable 13,795 Current liabilities Long-term debt Owners equity $136,720 $229,844 $390,574 Net fixed assets Total assets Total liabilities and owners equity GRAFFITI ADVERTISING, INC Balance Sheet as of December 31, 2016 $ 10,513 16,467 $ 26,980 $ 14,316 Accounts payable Cash Accounts receivable Inventory Current assets 21,098 Notes payable 22,755 Current liabilities 58,169 $152,100 $285,399 $464,479 Long-term debt Net fixed assets Total assets $406,310 Owners equity $464,479 ollabilities and owners equitya. Calculate the operating cash flow. (Dnot round intermediate calculations and round your answer to the nearest whole number, e.g., 32.) Operating cash flow b. Calculate the change in net working capital. (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.) Change in net working capta c. Calculate the net capital spending. (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.) Net capital spending d. Calculate the cash flow from assets. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.) Cash flow from assets e. Calculate the cash flow to creditors. (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.) Cash flow to creditors f. Calculate the cash flow to stockholders. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.) Cash flow to stockholders $
Answer #1

a) Cash Flow from Operations = EBIT - Taxes + Depreciation = 112373 - 37222 + 54,575 = 129726

b) Change in Net Working Capital = (Net working capital in 2016 - Net working capital in 2015) = (58169- 26980) - (46118 - 24010) =9081

c) Net capital spending = Total Long term debt in 2016 - Long term debt in 2015 = 152100 - 136720 = 15380

d) Cash flow From assets = Cash flow From Operations - Change in NWC - Change in Capital Spending = 129726 - 9081
- 15380 = 105,265